# Off-policy Evaluation with Deeply-abstracted States

### v2

@@ Appendix D Technical proofs
We provide the detailed proofs of our theorems (Lemma 1, Theorems 1 & 2) in this section.
Notations. For events or random variables $A,B,C$, $A\perp\!\!\!\perp B$ means the independence between $A$ and $B$ whereas $A\perp\!\!\!\perp B|C$ means the conditional independence between $A$ and $B$ given $C$.
An auxiliary lemma. To begin with, we introduce the following lemma which validates the unbiasedness of various OPE estimators under the model-free irrelevance conditions in Definitions 57, whose proof is given in Section D.1.
Lemma D.1 (Unbiasedness under model-free irrelevance conditions)
Lemma D.1 proves the unbiasedness of the four OPE methods presented in Section 3.3 when applied to the abstract state space, under the corresponding irrelevance conditions. Notably, DRL requires weaker irrelevance conditions compared to the Q-function-based method and MIS, owing to its inherent double robustness property.
D.1 Proof of Lemma D.1
We prove Lemma D.1 in this subsection. We first prove that under $Q^{\pi}$-, $\rho^{\pi}$- or $w^{\pi}$-irrelevance, the corresponding methods remain unbiased when applied to the abstract state space:
    $=$ $\sum_{a,s}\pi(a|s)Q_{\phi}^{\pi}(a,\phi(s))\mathbb{P}(S_{1}=s)$
    $=$ $\mathbb{E}[f_{1}(Q^{\pi}_{\phi})].$
    $\rho^{\pi}(a,s)=\sum_{s^{\prime}\in\phi^{-1}(x)}\mathbb{P}(S_{t}=s^{\prime}|A_{t}=a,\phi(S_{t})=x)\rho^{\pi}(a,s^{\prime}).$ (D.1)
    Notice that
    The denominator equals $b_{\phi,t}(a|x)$, the behavior policy when restricted to the abstract state space at time $t$. Notice that this behavior policy can be non-stationary over time, despite that $b$ being time-invariant. As for the numerator, it is straightforward to show that it equals $b(a|s^{\prime})\mathbb{P}(S_{t}=s^{\prime}|\phi(S_{t})=x)$. This together with (D.1) yields
    $\rho^{\pi}(a,s)=\sum_{s^{\prime}\in\phi^{-1}(x)}\frac{\pi(a|s^{\prime})}{b_{\phi,t}(a|x)}\mathbb{P}(S_{t}=s^{\prime}|\phi(S_{t})=x)=\frac{\pi_{\phi,t}(a|x)}{b_{\phi,t}(a|x)},$ (D.2)
    where $\pi_{\phi,t}$ denotes the target policy confined on the abstract state space at time $t$, namely, $\pi_{\phi,t}(a|x)=\sum_{s^{\prime}\in\phi^{-1}(x)}\pi(a|s^{\prime})\mathbb{P}(S_{t}=s^{\prime}|\phi(S_{t})=x)$. The last term in (D.2) is given by $\rho_{\phi,t}^{\pi}$. Consequently, the sequential IS ratio $\rho_{1:t}^{\pi}$ is equal to $\prod_{k=1}^{t}\rho_{\phi,k}^{\pi}(A_{k},\phi(S_{k}))$. This in turn yields $\mathbb{E}[f_{2}(\rho^{\pi})]=\mathbb{E}[f_{2}(\rho^{\pi}_{\phi})]$.
    Similar to (D.2), under $w^{\pi}$-irreleavance, it follows that
D.2 Proof of Lemma 1
We prove Lemma 1 in this subsection. We first show $Q^{\pi}$-irrelevance under model-irrelevance &\&& $\pi$-irrelevance, and prove the Fisher consistency of the Q-function-based method. Then, we establish Fisher consistency of MIS. Next, we derive Fisher consistency of SIS. Finally, we prove the Fisher consistency of DRL.
    $Q^{\pi}(a,s^{(1)})=Q^{\pi}(a,s^{(2)}),$ (D.3)
    whenever $s^{(1)}$ and $s^{(2)}$ satisfy $\phi(s^{(1)})=\phi(s^{(2)})$. This demonstrates $Q^{\pi}$-irrelevance, which further implies $\mathbb{E}[f_{1}(Q^{\pi})]=\mathbb{E}[f_{1}(Q^{\pi}_{\phi})]$ according to Lemma D.1. We next establish the identifiability of $Q^{\pi}_{\phi}$.
    Define
    Under reward-irrelevance, we have
    $=$ $\mathcal{R}(a,s^{(1)})$
    $=$ $\mathcal{R}(a,s^{(2)})$
    $=$ $Q_{1}^{\pi}(a,s^{(2)}).$
    Together with $\pi$-irrelevance, we obtain that
    $=$ $\sum_{a\in\mathcal{A}}Q_{1}^{\pi}(a,s^{(2)})\pi(a|s^{(2)})$
    $=$ $V_{1}^{\pi}(s^{(2)}).$
    Suppose we have shown that the following holds for any $j<T$,
    $Q_{j}^{\pi}(a,s^{(1)})=Q_{j}^{\pi}(a,s^{(2)})\,\,\mbox{and}\,\,V_{j}^{\pi}(s^{(1)})=V_{j}^{\pi}(s^{(2)}),$ (D.4)
    whenever $\phi(s^{(1)})=\phi(s^{(2)})$. Our goal is to show (D.4) holds with $j=T$.
    We similarly define $Q_{j,\phi}^{\pi}$ and $V_{j,\phi}^{\pi}$ as the Q- and value functions on the abstract state space. Similar to the proof of Theorem D.1, we can show that $Q_{j}^{\pi}=Q_{j,\phi}^{\pi}$ and $V_{j}^{\pi}=V_{j,\phi}^{\pi}$ for any $j<T$.

### v3

@@ Appendix D Technical proofs
We provide the detailed proofs of our theorems (Theorems 1 & 2 and Lemma 2) in this section.
Notations. For events or random variables $A,B,C$, $A\perp\!\!\!\perp B$ means the independence between $A$ and $B$ whereas $A\perp\!\!\!\perp B|C$ means the conditional independence between $A$ and $B$ given $C$. Let $c$ be a positive constant, which may take different values in different contexts. For two positive sequences $\{a_{n}\}_{n\geq 1}$, $\{b_{n}\}_{n\geq 1}$, we write $a_{n}=O(b_{n})$ if there exists a positive constant $c$ such that $a_{n}\leq c\cdot b_{n}$, and we write $a_{n}=o(b_{n})$ if $a_{n}/b_{n}\rightarrow 0$. Furthermore, if $a_{n}=O(b_{n})$ is satisfied, we write $a_{n}\lesssim b_{n}$. If $a_{n}\lesssim b_{n}$ and $b_{n}\lesssim a_{n}$, we write it as $a_{n}\asymp b_{n}$ for short. For any function $f$, denote $\|f\|=\sqrt{\mathbb{E}[f(X)^{2}]}$ with $X$ being random vector, $\mathbb{P}_{n}f=(NT)^{-1}\sum_{i=1}^{N}\sum_{j=1}^{T}f(A_{i,j},S_{i,j})$,
$\widehat{\Sigma}=\mathbb{P}_{n}\left[-\gamma\sum_{a^{\prime}}\pi(a^{\prime}|S^{\prime})\varphi\left(a^{\prime},S^{\prime}\right)\varphi(A,S)^{\top}+\varphi(A,S)\varphi(A,S)^{\top}\right],$
with $\Sigma=\mathbb{E}\left[-\gamma\sum_{a}\pi(a|S^{\prime})\varphi\left(a,S^{\prime}\right)\varphi(A,S)^{\top}+\varphi(A,S)\varphi(A,S)^{\top}\right]$ being the population counterpart.
Auxiliary lemmas. To begin with, we introduce the subsequent lemmas. One lemma exhibits the bias-variance decomposition of diverse OPE estimators for MDP, whose proof is provided in Section D.2. Additionally, we present the properties regarding the linear estimation of $Q$ and $w$, with the corresponding proof furnished in Section D.3.
Lemma D.1 (Bias-variance decomposition for the MSE)
Assume Assumptions 14 hold. Suppose that $\|\widehat{Q}^{\pi}-{Q}^{\pi}\|=O(R_{\mathrm{max}}(1-\gamma)^{-1}\kappa_{q})$, and $\|\widehat{w}^{\pi}-{w}^{\pi}\|=O(\kappa_{w})$ are satisfied.
  • The MSE of DRL (Method 4) applied to the MDP is given by:
    $\frac{1}{N}\mathbb{V}\left[V^{\pi}\left(S\right)\right]+\frac{1}{NT(1-\gamma)^{2}}\mathbb{V}\left[w^{\pi}(A,S)\left(R+\gamma V^{\pi}\left(S^{\prime}\right)-Q^{\pi}(A,S)\right)\right]$
    $+O\left(\frac{R_{\max}^{2}\max(\kappa_{q},\kappa_{w},N^{-1/2})\log(N)}{(1-\gamma)^{2}N}+\frac{R_{\max}^{2}}{\sqrt{N}(1-\gamma)^{2}}\max\{\kappa_{q}\kappa_{w},\sqrt{N}\kappa_{q}^{2}\kappa_{w}^{2}\}\right).$ (D.1)
  • When both the Q-function and MIS ratio are estimated via linear function approximation, the MSEs of Q-function-based estimator (Method 1) and MIS (Method 3) applied to the MDP identical to that of DRL.
  • When $\rho^{\pi}$ is known, the MSE of SIS (Method 2) applied to the MDP is upper bounded by
    $\frac{4R_{\max}^{2}}{NT^{2}(1-\gamma)^{2}}\sum_{t=1}^{T}\gamma^{2t-2}\mathbb{E}(\rho_{1:t}^{\pi})^{2}+\frac{\gamma^{2T}R_{\max}^{2}}{(1-\gamma)^{2}}.$
Before we figure out the value of $\kappa_{q}$ and $\kappa_{w}$, we need the following assumption.
Assumption 5 (Positive definite)
(i) $\lambda_{\min}(\Sigma)\geq{c}$ for some constant ${c}>0$ where $\lambda_{\min}(K)$ denotes the minimum eigen value of a matrix $K$.
Lemma D.2 (The order of $\widehat{Q}^{\pi}$ and $\widehat{w}^{\pi}$)
Suppose Assumptions 15 hold. When we utilize a linear model to estimate $Q$ and $w$, within the tabular setting, it can be demonstrated that both $\kappa_{q}$ and $\kappa_{w}$ are of the order $\kappa_{q}$ and $\kappa_{w}$ are $(NT)^{-1/2}|\mathcal{S}||\mathcal{A}|$ and $N^{-1/2}|\mathcal{S}||\mathcal{A}|$, respectively.
Based on Lemmas D.1 and D.2, we can get the bias-variance decomposition for the MSE under tabular setting.
Corollary 1 (Bias-variance decomposition for MSE in tabular setting)
Assume Assumptions 15 hold. The bias-variance decomposition MSE of DRL is given by:
$\frac{1}{N}\mathbb{V}\left[V^{\pi}\left(S\right)\right]+\frac{1}{NT(1-\gamma)^{2}}\mathbb{V}\left[w^{\pi}(A,S)\left(R+\gamma V^{\pi}\left(S^{\prime}\right)-Q^{\pi}(A,S)\right)\right]$
$+O\left(\frac{R_{\max}^{2}|\mathcal{S}||\mathcal{A}|\log(N)}{(1-\gamma)^{2}N^{3/2}}\right)+O(\frac{R_{\max}^{2}}{\sqrt{N}(1-\gamma)^{2}}\frac{|\mathcal{S}|^{2}|\mathcal{A}|^{2}}{NT^{1/2}}).$
D.1 Proof of Theorem 1
Since $\phi$ is MSA, then Theorem 1 directly follows from Lemma D.1.
D.2 Poof of Lemma D.1
In the following, we calculate the MSE of DRL, Q-function-based method, MIS and SIS one by one.
  • MSE of DRL. Our DRL estimator based on the dataset $\mathcal{D}$ is
    $\widehat{J}_{\mathrm{DRL}}\left(\widehat{Q}^{\pi},\widehat{w}^{\pi}\right)=$ $\mathbb{P}_{n}\left[\psi\left(\widehat{Q}^{\pi},\widehat{w}^{\pi}\right)\right]$
    $=$ $\frac{1}{N}\sum_{i=1}^{N}\pi(A_{i,1}|S_{i,1})\widehat{Q}^{\pi}\left(A_{i,1},S_{i,1}\right)+(1-\gamma)^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\widehat{w}^{\pi}\left(A_{i,j},S_{i,j}\right)$
    ×$\times\left(R_{i,j}+\gamma\sum_{a}\pi(a|S_{i,j+1})\widehat{Q}^{\pi}\left(a,S_{i,j+1}\right)-\widehat{Q}^{\pi}\left(A_{i,j},S_{i,j}\right)\right),$
    where $\widehat{Q}^{\pi},\widehat{w}^{\pi}$ are some plug-in estimates of $Q^{\pi}$ and $w^{\pi}$, respectively. Now, we will calculate the MSE of $\widehat{J}_{\mathrm{DRL}}$. A direct calculation yields
    $\mathbb{P}_{n}\psi\left[\left(\widehat{Q}^{\pi},\widehat{w}^{\pi}\right)\right]-J(\pi)$
    $=$ $\left\{\mathbb{P}_{n}\left[\psi\left(\widehat{Q}^{\pi},\widehat{w}^{\pi}\right)-\psi\left(w^{\pi},Q^{\pi}\right)\right]+\left\{\mathbb{P}_{n}\left[\psi\left(w^{\pi},Q^{\pi}\right)\right]-J(\pi)\right\}\right\}$
    $=$ $\mathbb{P}_{n}\left[\psi\left(\widehat{Q}^{\pi},\widehat{w}^{\pi}\right)-\psi\left(w^{\pi},Q^{\pi}\right)\right]+\left\{\mathbb{P}_{n}\left[\psi\left(w^{\pi},Q^{\pi}\right)\right]-J(\pi)\right\}$
    $=$ $I_{1}+I_{2}.$
    For $I_{1}$, we have
    $I_{1}=$ $\mathbb{P}_{n}\left[\psi\left(\widehat{Q}^{\pi},\widehat{w}^{\pi}\right)-\psi\left(w^{\pi},Q^{\pi}\right)\right]$
    $=$ $\mathbb{P}_{n}\left[\psi\left(\widehat{Q}^{\pi},\widehat{w}^{\pi}\right)-\psi\left({w}^{\pi},\widehat{Q}^{\pi}\right)\right]$
    $+\mathbb{P}_{n}\left[\psi\left({w}^{\pi},\widehat{Q}^{\pi}\right)-\psi\left(w^{\pi},Q^{\pi}\right)\right]$
    $=$ $\frac{1}{N}\sum_{i=1}^{N}\pi(A_{i,1}|S_{i,1})\left\{\widehat{Q}^{\pi}\left(A_{i,1},S_{i,1}\right)-{Q}^{\pi}\left(A_{i,1},S_{i,1}\right)\right\}$
    $+(1-\gamma)^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}{w}^{\pi}\left(A_{i,j},S_{i,j}\right)\left(\gamma\sum_{a}\pi(a|S_{i,j+1})\left\{\widehat{Q}^{\pi}\left(a,S_{i,j+1}\right)-{Q}^{\pi}\left(A_{i,j+1},S_{i,j+1}\right)\right\}\right.$
    $\left.-\left\{\widehat{Q}^{\pi}\left(A_{i,j},S_{i,j}\right)-{Q}^{\pi}\left(A_{i,j},S_{i,j}\right)\right\}\right)+(1-\gamma)^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\left\{\widehat{w}^{\pi}\left(A_{i,j},S_{i,j}\right)-{w}^{\pi}\left(A_{i,j},S_{i,j}\right)\right\}$
    ×$\times\left(\gamma\sum_{a}\pi(a|S_{i,j+1})\left\{\widehat{Q}^{\pi}\left(a,S_{i,j+1}\right)-{Q}^{\pi}\left(a,S_{i,j+1}\right)\right\}-\left\{\widehat{Q}^{\pi}\left(A_{i,j},S_{i,j}\right)-{Q}^{\pi}\left(A_{i,j},S_{i,j}\right)\right\}\right)$
    $+(1-\gamma)^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\left\{\widehat{w}^{\pi}\left(A_{i,j},S_{i,j}\right)-{w}^{\pi}\left(A_{i,j},S_{i,j}\right)\right\}$
    ×$\times\left\{R_{i,j}+\gamma\sum_{a}\pi(a|S_{i,j+1}){Q}^{\pi}\left(a,S_{i,j+1}\right)-{Q}^{\pi}\left(A_{i,j},S_{i,j}\right)\right\}.$ (D.2)
    Denote
    $I_{11}=$ $\frac{1}{N}\sum_{i=1}^{N}\pi(A_{i,1}|S_{i,1})\left\{\widehat{Q}^{\pi}\left(A_{i,1},S_{i,1}\right)-{Q}^{\pi}\left(A_{i,1},S_{i,1}\right)\right\}+(1-\gamma)^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}{w}^{\pi}\left(A_{i,j},S_{i,j}\right)$
    ×$\times\left(\gamma\sum_{a}\pi(a|S_{i,j+1})\left\{\widehat{Q}^{\pi}\left(a,S_{i,j+1}\right)-{Q}^{\pi}\left(a,S_{i,j+1}\right)\right\}-\left\{\widehat{Q}^{\pi}\left(A_{i,j},S_{i,j}\right)-{Q}^{\pi}\left(A_{i,j},S_{i,j}\right)\right\}\right),$ (D.3)
    $I_{12}=(1-\gamma)^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\left\{\widehat{w}^{\pi}\left(A_{i,j},S_{i,j}\right)-{w}^{\pi}\left(A_{i,j},S_{i,j}\right)\right\}$
    ×$\times\left(\gamma\sum_{a}\pi(a|S_{i,j+1})\left\{\widehat{Q}^{\pi}\left(a,S_{i,j+1}\right)-{Q}^{\pi}\left(a,S_{i,j+1}\right)\right\}-\left\{\widehat{Q}^{\pi}\left(A_{i,j},s_{i,j}\right)-{Q}^{\pi}\left(A_{i,j},S_{i,j}\right)\right\}\right),$ (D.4)
    and
    $I_{13}=$ $(1-\gamma)^{-1}\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\left\{\widehat{w}^{\pi}\left(A_{i,j},S_{i,j}\right)-{w}^{\pi}\left(A_{i,j},S_{i,j}\right)\right\}$
    ×$\times\left\{R_{i,j}+\gamma\sum_{a}\pi(a|S_{i,j+1}){Q}^{\pi}\left(a,S_{i,j+1}\right)-{Q}^{\pi}\left(A_{i,j},S_{i,j}\right)\right\}.$ (D.5)
    It follows from Assumption 4 that we allow the function classes to be infinite hypothesis classes. To handle this, we can find $\epsilon$-nets of $\mathcal{Q}$ and $\mathcal{W}$, denoted by $\mathcal{Q}_{0}$ and $\mathcal{W}_{0}$ with $\epsilon$ proportional to $N^{-1}R_{\max}/(1-\gamma)$ and $N^{-1}$, respectively. This kind of restricting to the finite hypothesis class $\mathcal{Q}_{0}$ and $\mathcal{W}_{0}$ provides reasonable approximations for $\mathcal{Q}$ and $\mathcal{W}$ with the approximation error upper bounded by
    $O\left(\frac{R_{\max}}{(1-\gamma)N}\right)\,\,\mathrm{and}\,\,O\left(\frac{1}{N}\right).$
    Meanwhile, the number of elements in $\mathcal{Q}_{0}$ and $\mathcal{W}_{0}$ is of the order $O\left(N^{V_{q}}\right)$ and $O\left(N^{V_{w}}\right)$ for some positive constants $V_{q}$ and $V_{w}$.
    Now, we analyze the error bound of $I_{11}$. Fix $f\in\mathcal{Q}_{0}$ and define
    $X_{i}\left(f\right)=$ $\pi(A_{i,1}|S_{i,1})\left\{f\left(A_{i,1},S_{i,1}\right)-Q^{\pi}\left(A_{i,1},S_{i,1}\right)\right\}+(1-\gamma)^{-1}\frac{1}{T}\sum_{j=1}^{T}{w}^{\pi}\left(A_{i,j},S_{i,j}\right)$
    ×$\times\left(\gamma\sum_{a}\pi(a|S_{i,j+1})\left\{f\left(a,S_{i,j+1}\right)-Q^{\pi}\left(a,S_{i,j+1}\right)\right\}-\left\{f\left(A_{i,j},S_{i,j}\right)-Q^{\pi}\left(A_{i,j},S_{i,j}\right)\right\}\right),$ (D.6)
    It is easy to see
    $I_{11}=\frac{1}{N}\sum_{i=1}^{N}X_{i}\left(\widehat{Q}^{\pi}\right)\,\,\mathrm{and}\,\,\mathbb{E}[X(\widehat{Q}^{\pi})]=0.$
    Since $X_{i}\left(f\right)$ are i.i.d., we apply Bernstein’s inequality and the union bound over all $f\in\mathcal{Q}_{0}$. With probability at least $1-\delta$, we have
    $|\frac{1}{n}\sum_{i=1}^{n}X_{i}\left(f\right)-\mathbb{E}\left[X\left(f\right)\right]|$
    $\leq$ $\sqrt{\frac{2\mathbb{V}\left[X\left(f\right)\right]\ln\frac{2|\mathcal{Q}_{0}|}{\delta}}{N}}+\frac{2cR_{\max}\ln\frac{2|\mathcal{Q}_{0}|}{\delta}}{3N(1-\gamma)}.$ (D.7)
    It follows from Assumption 4, and $\|\widehat{Q}^{\pi}(A,S)-{Q}^{\pi}(A,S)\|=O(R_{\mathrm{max}}(1-\gamma)^{-1}\kappa_{q})$ and Jensen’s inequality that
    $\|\sum_{a}\pi(a|S)\big{[}\widehat{Q}^{\pi}(a,S)-{Q}^{\pi}(a,S)\big{]}\|=O(R_{\mathrm{max}}(1-\gamma)^{-1}\kappa_{q}).$ (D.8)
    Combing Assumption 4 and equation (D.8), we can obtain
    $\mathbb{V}\left[X\left(\widehat{Q}^{\pi}\right)\right]\lesssim(1-\gamma)^{-2}R_{\max}^{2}\kappa_{q}^{2}.$ (D.9)
    More specifically, by substituting (D.9) into (D.2), it can be shown that with probability at least $1-O\left(N^{-2}\right)$, $I_{11}$ can be upper bounded by
    $O\left(\frac{R_{\max}\log(N)}{(1-\gamma)N}\right)+O\left(\frac{R_{\max}\kappa_{q}\sqrt{\log(N)}}{(1-\gamma)\sqrt{N}}\right).$ (D.10)
    By similar arguments as above, it can be shown that with probability at least $1-O\left(N^{-2}\right)$, $I_{13}$ can be upper bounded by
    $O\left(\frac{R_{\max}\log(N)}{(1-\gamma)N}\right)+O\left(\frac{R_{\max}\kappa_{w}\sqrt{\log(N)}}{(1-\gamma)\sqrt{N}}\right).$ (D.11)
    For $I_{12}$, direct calculations yield
    $\mathbb{E}|I_{12}|^{2}\lesssim(1-\gamma)^{-2}R_{\max}^{2}\kappa_{q}^{2}\kappa_{w}^{2}.$ (D.12)
    Thus, we can get
    $\mathbb{E}[I_{1}^{2}]\leq$ $O\left(\frac{R_{\max}^{2}\log^{2}(N)}{(1-\gamma)^{2}N^{2}}\right)+O\left(\frac{R_{\max}^{2}\kappa_{w}^{2}\log(N)}{(1-\gamma)^{2}N}\right)+O(\frac{R_{\max}^{2}}{(1-\gamma)^{2}}\kappa_{q}^{2}\kappa_{w}^{2})$
    $+O\left(\frac{R_{\max}^{2}\kappa_{q}^{2}\log(N)}{(1-\gamma)^{2}N}\right)+O(\frac{R_{\max}^{2}}{N^{2}(1-\gamma)^{2}}).$ (D.13)
    For $I_{2}$, direct calculations yield
    $\mathbb{E}[I_{2}^{2}]=\frac{1}{N}\mathbb{V}\left[V^{\pi}\left(S\right)\right]+\frac{1}{NT(1-\gamma)^{2}}\mathbb{V}\left[w^{\pi}(A,S)\left(R+\gamma V^{\pi}\left(S^{\prime}\right)-Q^{\pi}(A,S)\right)\right].$ (D.14)
    By Cauchy-Schwartz inequality, (D.2) and (D.14), we can get
    $\mathbb{E}[I_{1}I_{2}]\leq$ $\sqrt{\mathbb{E}[I_{1}^{2}]}\sqrt{\mathbb{E}[I_{2}^{2}]}$
    $\leq\frac{R_{\max}}{(1-\gamma)\sqrt{N}}\left\{O\left(\frac{R_{\max}\max(\kappa_{q},\kappa_{w},\sqrt{N^{-1}\log(N)})\sqrt{\log(N)}}{(1-\gamma)\sqrt{N}}\right)+O(\frac{R_{\max}}{(1-\gamma)}\kappa_{q}\kappa_{w})\right\}.$ (D.15)
    Combing (D.2), (D.14) and (D.2), we can get
    $\mathrm{MSE}_{\mathrm{DRL}}=$ $\mathbb{E}[\widehat{J}_{\mathrm{DRL}}\left(\widehat{Q}^{\pi},\widehat{w}\right)-J(\pi)]^{2}$
    $=$ $\frac{1}{N}\mathbb{V}\left[V^{\pi}\left(S\right)\right]+\frac{1}{NT(1-\gamma)^{2}}\mathbb{V}\left[w^{\pi}(A,S)\left(R+\gamma V^{\pi}\left(S^{\prime}\right)-Q^{\pi}(A,S)\right)\right]$
    $+O\left(\frac{R_{\max}^{2}\max(\kappa_{q},\kappa_{w},N^{-1/2})\log(N)}{(1-\gamma)^{2}N}+\frac{R_{\max}^{2}}{\sqrt{N}(1-\gamma)^{2}}\max\{\kappa_{q}\kappa_{w},\sqrt{N}\kappa_{q}^{2}\kappa_{w}^{2}\}\right).$
  • MSE of Q-function based method. It follows from Example 7 in Uehara et al. [2020] and $\widehat{\Sigma}$ is non-sigular that, when we use the linear method to estimate $Q^{\pi}$ by MQL, $\widehat{Q}^{\pi}(a,s)=\varphi(a,s)^{\top}\widehat{\beta}$ where $\phi(a,s)\in$ $\mathbb{R}^{d}$ is some basis function and $\widehat{\beta}$ takes the form
    $\widehat{\Sigma}^{-1}\mathbb{P}_{n}[R\varphi(A,S)].$
    A direct calculation yields
    $\widehat{\beta}^{\top}\mathbb{P}_{n}\left[-\gamma\sum_{a^{\prime}}\pi(a^{\prime}|S^{\prime})\varphi\left(a^{\prime},S^{\prime}\right)\varphi(A,S)^{\top}+\varphi(A,S)\varphi(A,S)^{\top}\right]-\mathbb{P}_{n}[R\varphi(A,S)^{\top}]=0.$
    Consider $w^{\pi}\in\mathcal{F}$ whose parameter is $\alpha$, and denote $\widehat{w}^{\pi}(a,s)=\varphi(a,s)^{\top}\widehat{\alpha}$. It is easy to verify
    $0=$ $\left\{\widehat{\beta}^{\top}\mathbb{P}_{n}\left[-\gamma\sum_{a^{\prime}}\pi(a^{\prime}|S^{\prime})\varphi\left(a^{\prime},S^{\prime}\right)\varphi(A,S)^{\top}+\varphi(A,S)\varphi(A,S)^{\top}\right]-\mathbb{P}_{n}[R\varphi(A,S)^{\top}]\right\}\widehat{\alpha}$
    $=$ $\mathbb{P}_{n}\left[-\gamma\sum_{a^{\prime}}\pi(a^{\prime}|S^{\prime})\widehat{\beta}^{\top}\varphi\left(a^{\prime},S^{\prime}\right)\varphi(A,S)^{\top}\widehat{\alpha}+\widehat{\beta}^{\top}\varphi(A,S)\varphi(A,S)^{\top}\widehat{\alpha}\right]-\mathbb{P}_{n}[R\varphi(A,S)^{\top}\widehat{\alpha}]$
    $=$ $-\frac{1}{NT}\sum_{i=1}^{N}\sum_{j=1}^{T}\widehat{w}^{\pi}\left(A_{i,j},S_{i,j}\right)\left(R_{i,j}+\gamma\sum_{a}\pi(a|S_{i,j+1})\widehat{Q}^{\pi}\left(a,S_{i,j+1}\right)-\widehat{Q}^{\pi}\left(A_{i,j},S_{i,j}\right)\right).$
    This means $\widehat{J}_{\mathrm{Q}}(\widehat{Q}^{\pi})=\widehat{J}_{\mathrm{DLR}}(\widehat{Q}^{\pi},\widehat{w}^{\pi})$. Thus, we have
    $\mathrm{MSE}_{\mathrm{Q}}=\mathrm{MSE}_{\mathrm{DRL}}.$
  • MSE of MIS. We assume that both $\mathcal{W}$ and $\mathcal{Q}$ are set to the same linear function space =$\mathcal{F}=\{(a,s)$ $\left.\varphi(a,s)^{\top}\beta:\beta\in\mathbb{R}^{d}\right\}$, and $\widehat{\Sigma}$ is non-singular. It follows from Example 2 in Uehara et al. [2020] that, the estimator of $w^{\pi}(a,s)$ by MWL takes the following closed form:
    $\widehat{w}^{\pi}(a,s)=\varphi(a,s)^{\top}\widehat{\alpha},$
    with
    $\widehat{\alpha}=$ $\widehat{\Sigma}^{-1}(1-\gamma)\frac{1}{N}\sum_{i=1}^{N}\left[\sum_{a}\pi(a|S_{i,1})\varphi\left(a,S_{i,1}\right)\right].$